Find the largest possible area of a shaded rectangle which lies within parabola "g(x) = - x^2 + 12" which has its base on the x-axis and two of its vertices lying on the parabola and above the x-axis.


1
Expert's answer
2021-05-19T18:09:16-0400

Question:-

Find the largest possible area of a shaded rectangle which lies within parabola "g(x) = - x^2 + 12" which has its base on the x-axis and two of its vertices lying on the parabola and above the x-axis.


Solution:-

according to question,

we get diagram,

such that as image




here,

coordinate of A and B ,assume (-x,0) and (x,0) respectively.

and

coordinate of C and D will be

(x,(-x2+12)) and (-x,(-x2+12)) respectively, becouse y coordinate given as (-x2+12) in the question.


now we find largest possible area of rectangle.

here in image,


rectangle lenght=AB=2x

rectangle width=CD=(-x2+12)


So

rectangle area=2x.(-x2+12)

let it is a function of x, such that

f(x)=2x.(-x2+12)


now we want largest value of this,

so we will maxima this,


so we find f'(x),

f'(x)=2(-x2+12)+2x(-2x)


now put f'(x)=0

2(-x2+12)+2x(-2x)=0

-6x2+24=0

6x2=24

x=±2


now put positive value of x in the f(x)

and we will be get largest area of rectengle.


so put

f(2)=2(2)(-22+12)

=32


largest possible area of a rectangle= 32